OpenStudy (anonymous):

Explain, in complete sentences, why the expression x^2 + 16x – 48 is prime?

OpenStudy (anonymous):

The discriminant of x2 + 16x – 48 is 16^2 - (4)(1)(-48) = 256 + 192 = 448 . Since 448 is not a perfect square, the roots (and factors) will be irrational.

OpenStudy (anonymous):

thanks again ;)